# Motion simulation of radio-labeled cells in whole-body positron emission tomography

### v2

@@ Abstract
Cell tracking is a subject of active research gathering great interest in medicine and biology. Positron emission tomography (PET) is well suited for tracking radio-labeled cells in vivo due to its exceptional sensitivity and whole-body capability. For validation, ground-truth data is desirable that realistically mimics the flow of cells in a clinical situation. This study develops a workflow (CeFloPS) for simulating moving radio-labeled cells in a human phantom. From the XCAT phantom, the blood vessels are reduced to nodal networks along which cells can move and distribute to organs and tissues. The movement is directed by the blood flow which is calculated in each node using the Hagen-Poiseuille equation and Kirchhoff’s laws assuming laminar flow. Organs are voxelized and movement of cells from artery entry to vein exit is generated via a biased 3D random walk. The probabilities of whether cells move or stay in tissues are derived from rate constants of physiologically based compartment modeling. PET listmode data is generated using the Monte-Carlo simulation framework GATE based on the definition of a large-body PET scanner with cell paths as moving radioactive sources and the XCAT phantom providing attenuation data. From the flow simulation of 100,000 cells, 100 sample cells were further processed by GATE and listmode data was reconstructed into images for comparison. As demonstrated by comparisons of simulated and reconstructed cell distributions, CeFloPS has the capacity to simulate the cell behavior of whole-body PET. It achieves this simulation in a way that is anatomically and physiologically reasonable, thereby providing valuable data for the development and validation of cell tracking algorithms.
@@ 1 Introduction
Cell imaging and tracking is a subject of active research gathering great interest in the medical and biomedical science community. Non-invasive imaging techniques using fluorescent or luminescent markers in optical imaging (OI), iron oxide particles in magnetic resonance imaging (MRI), or radio-labeled tracers in single photon emission computed tomography (SPECT) and positron emission tomography (PET) have been frequently used in preclinical studies to permit in vivo visualization of living cells [1]. For human application, MRI, PET and SPECT are the most promising tools as they provide either high spatial resolution (MRI) or excellent whole-body sensitivity (PET, SPECT). Of these imaging modalities, PET has become of special interest following labeled immune cells in the context of stem cell imaging and cell-based therapies [2].
PET is particularly advantageous because it combines exceptional sensitivity with the possibility of whole-body coverage which has been further advanced in recent years by the design of large-field-of-view (LAFOV) PET systems with extended coverage of the axial field-of-view in the range of 1 to 2 meters, improved time-of-flight capabilities and excellent system sensitivities [3]. Using either long-lived isotopes or reporter genes, immunoPET has great potential to identify biomarkers allowing to treat patients based on their most likely response to a dedicated immunotherapy [4, 5, 6].
The idea to image or track biological cells with PET is highly related to the field of Positron Emission Particle Tracking (PEPT) which has been developed primarily for technical applications, such as evaluation of particle–fluid or particle–particle interactions in the pharmaceutical, chemical, oil, mining, construction and power generation industries [7]. The idea is to follow one or a few radio-labeled particles by tracking the coincidence lines-of-response (LORs) recognized from its two back-to-back gamma rays emitted after beta-plus decay. Thereby, a trajectory can be calculated by connecting LORs in a most probable way. As PEPT applications are generally carried out outside a clinical environment, special industrially usable devices were developed in parallel to the invention of clinical PET systems [8, 9]. Following the idea of PEPT in a clinical or pre-clinical environment, the outstanding sensitivity of PET can be utilized to track radio-labeled singular cells traveling inside a biological system (e.g. human, mouse) [9, 10, 11, 12].
@@
The second simulation enables to generate raw PET data from the simulated cell pathways using Monte Carlo based physics simulations. Such simulations were also used in the PEPT context based on the Geant4 Application for Tomographic Emission (Gate) [18]. E.g. Herald et al. simulated both the ADAC/Phillips Forte and the Siemens Inveon pre-clinical PET and compared fluid simulations with experimental data [19, 20]. The foundation of our Gate simulation, developed here, is the concept of new total-body PET with an extensive field of view and high sensitivity, which presents novel possibilities for human whole-body cell tracking.
The task of cell tracking is not part of this work, but has been developed in many variants by different groups in the past [21, 5]. Nguyen et al. recently presented a new algorithm, PEPT-EM, which was able to successfully track the fate of more than 70 radio-labeled melanoma cells in a mouse model [12]. In this context, the group recently introduced an image reconstruction algorithm for PET that employs an optimal transport-based regularisation technique. This algorithm demonstrates considerable potential for the tracking of radioactively labelled cells within a human body, with no limitation on the number of trackable cells, and with the capacity to address the inherent handling of scattered, random and background events. The proposed methodology integrates the information from all detected events to facilitate the reconstruction of dynamic evolution, whilst concurrently enhancing the reconstruction quality through the enforcement of temporal consistency. [22, 23, 24].
@@ 2.1 Simulation basis
CeFloPS is based on anatomical data provided by the XCAT phantom with high resolution 3D models of all organs, blood vessels, bones and muscles (Fig. 1a) in the form of STEP files [25]. Using Inventor (Version 2023, Autodesk Inc., San Rafael, USA), these STEP files were converted into StL files (StereoLithography file format), providing triangulated surface meshes of all organs and structures. The StL files were further processed in Python (Version 3.11.5, Python Software Foundation) with the help of the trimesh library [26], allowing for operations on the meshes and giving access to the coordinates of the vertices. For further processing, meshes of all muscles and internal organs were converted into voxelized volumes with voxel size 0.5 mm $\times$ 0.5 mm $\times$ 0.5 mm, while the complex anatomy of blood vessel meshes was simplified by reducing it to node networks. The nodes describe the possible paths a cell can take and are placed along the central axis of a blood vessel, storing information about (i) its spatial coordinates, (ii) which nodes are adjacent in the direction of blood flow and (iii) the radius of the blood vessel.
To generate the nodes, the blood vessels were divided into segments starting from one end point of a vessel. For each segment, the barycenter was formed of all vertices located in this segment, resulting in points lying inside a blood vessel. Starting from the first point inside a blood vessel, planes, orthogonal to the blood vessel, were then placed at an interval of 1 mm along the vectors between these points. Using the intersection of those planes with the blood vessel surface, the vessel radius and the new center point were calculated, resulting in the final 1 mm spaced nodes positions (see Fig. 2).
Nodes that lie at bifurcations of blood vessels and thus have more than one neighbor in the direction of blood flow were marked and are in the following referred to as links. Since the spatial resolution of the XCAT phantom is limited, some connections between vessels are missing, as seen in Fig. 1b. In this case we interpolate linearly between the nodes of those vessels to generate a connection. In addition, the spatial resolution of blood vessels varies from organ to organ. The smallest radius defined in XCAT is 0.15 mm and is located in the toes. Consequently, no capillary bed is included that connects the nodes of arteries and veins. (Fig. 1c). This results in a total of five separate networks of nodes: arteries, veins, portal vein, and pulmonary arteries and veins, all of which, except for the portal vein, are connected via the heart. In order to compensate for the lack of capillaries, these separate networks are artificially combined to form a single closed network in which opposing arteries and veins leading to the same organ are virtually connected.
@@ 2.2 Blood flow simulation
From the node networks, the mean blood flow rate $\overline{Q}_{i}$ as well as the vessel resistance $R_{i}$ were estimated for each node $i$ from a highly simplified flow model using Kirchhoff’s laws in combination with the Hagen-Poiseuille equation. This was done by assuming laminar flow in every vessel, treated as a cylinder [27, 28]:
@@
The individual velocity of a single cell is assigned by pseudo-random number sampling based on a cumulative distribution function (CDF). This ensures that a group of cells starting from the same node have different velocities and disperse over time in the blood vessels. The CDF was derived using the velocity profile of laminar flow in a tube [29]
@@ 2.3 Organ distribution simulation
Inside a voxelized organ, cell motion is guided by a biased 3D random walk along the voxels to simulate the capillary bed and the tissue exchange. This process is carried out in a series of discrete calculations, each with a defined time step. The current simulations have set the time step to 600 ms. To commence, voxels containing an artery entrance or a vein exit were identified and subsequently referred to as start-voxels and end-voxels, respectively. If an organ does not have an artery entrance and/or vein exit due to missing connections in the XCAT phantom, the nearest artery/vein is calculated and the closest voxel to that blood vessel is determined as start-/end-voxel.
Remaining blood vessels are connected to an organ in the same way if their distance to that organ lies within a 2 cm threshold. To compensate for missing arterial branches in muscles and organs, each node of the arteries and veins can serve as entrance and exit for cells, respectively. For every voxel within an organ or muscle, direction vectors towards all start- and end-voxels were determined along with separate weighting values. As the distance $d$ between a voxel and its start/end voxel increases, these weightings decrease by a factor of $1/d^{2}$, as shown in Fig. 3. The direction of the vectors is pointing away from start-voxel or towards end-voxel. The combination of those vectors represents the most probable directions for a cell to move. The weightings of vectors in start- and end-voxels were set to zero.
The random walk is initiated at a start-voxel, where a cell can move to one of the surrounding voxels. This gives 26 possible directions (provided the start-voxel does not reside on the organ’s periphery), all with the same probability. For the proceeding steps, the direction is biased in accordance with the previously determined weightings of the direction vectors in each voxel. If a cell reaches an end-voxel, the random walk concludes, followed by the cell exiting via the vein back into the nodal network for further blood flow simulation. The speed at which cells perform the random walk in the different organs is calibrated using the delay times determined by [30]. For this purpose, first the distribution of required time to travel from start- to end-voxel in each organ is determined for cells at an arbitrary random walk speed. The mean value of this distribution is calculated and compared with the value derived by [30], leading to new speed values for the corresponding organ.
@@
In addition to the biased random walk, a compartment model was incorporated into the organ distribution simulation to account for the transition state of a cell in tissue. Besides being spatially located in a voxel, a cell is also in a compartment state, which dictates whether a cell can move or not. The exchange between compartments is determined by transition rates $k_{i}$. The number of compartments and the values of the transition rates must be known in advance to the CeFloPS simulation and can be obtained from (physiologically based) kinetic models. As a first example and as proof-of-principle, we employ a two-compartment model comprising the compartments $C_{1}$ and $C_{2}$ (shown in Fig. 4). The compartment $C_{1}$ signifies a reversible binding of the cells, while $C_{2}$ denotes an irreversible binding. Consequently, if a cell is situated within either $C_{1}$ or $C_{2}$, it does not execute a random walk and remains temporarily or permanently in the same voxel. The change in amount of cells in these compartments can be described with the transition rates $K_{1}$, $k_{2}$, and $k_{3}$ through the differential equations
@@
The arterial input function (AIF), denoted by $C_{A}(t)$, serves as a metric for the concentration or number of cells present in the arterial blood and also has to be entered as input into the CeFloPS simulation. One method for generating an AIF is outlined in [31]. It is important to note that, in general, transition rates are only defined in the underlying kinetic model for a subset of the tissues included in CeFloPS. Consequently, the sum of all cells transitioning from the arterial blood into the specified tissues may not be equal to the rate of change of the AIF. In CeFlops, this would result in an erroneous distribution of cells, which would not correspond to (10). To counteract this, two virtual compartments were introduced. The first compartment, denoted by $C_{\mathrm{inj}}$, signifies the injection of cells into the arterial blood during the beginning of the simulation. The second compartment, designated by $C_{\mathrm{rem}}$, describes cells which are removed from the arterial blood and go into a tissue not defined in the kinetic model rates. The rate of change of the AIF is then given by
@@
The transition rates $k_{\mathrm{inj}}$ and $k_{\mathrm{rem}}$ are selected as time-dependent to account for the time-limited injection phase of the cells at the beginning of a simulation and potential delays during the transition into tissues. The determination of $k_{\mathrm{inj}}(t)$ and $k_{\mathrm{rem}}(t)$ is achieved through the implementation of a 3-layer neural network. The neural network identifies the optimal values for $k_{\mathrm{inj}}$ and $k_{\mathrm{rem}}$ at each simulation time $t$ by fitting the cell distributions in all tissues to fulfill the differential equations from (10) and (11). The latter phase of the distributions, in which a state of equilibrium is ultimately achieved, is given greater weight relative to the initial phase, during which significant alterations may occur. This ensures that the equilibrium phase at late stages of the simulation is mainly determined by the compartment model. The neural network is set up using the Python library PyTorch [32].
@@
In this example, we implement the rate constants of a 18F-Fluordesoxyglucose (FDG) model taken from [37] and listed in Table LABEL:tab:k_values. The application of kinetic modeling with FDG has been comprehensively addressed in the extant literature, thereby facilitating a more robust validation of our simulation compared to kinetic modeling with cells, which is still scarce in the literature and usually only takes few tissues into account. The simulated cells in this example therefore do not behave like real, existing cells, but like an FDG tracer. In this context, $C_{1}$ and $C_{2}$ denote the unbound and bound states of an FDG ”cell” in tissue and $C_{A}$ the free FDG ”cell” in the blood supply of the tissue. Based on the probabilities calculated with (12) from these transition rates, a random decision is made at each time step as to whether a cell transitions between these compartments. Fig. 5 demonstrates the temporal cell distributions calculated by the Markov chain approach in comparison to the whole population based compartment modelling using differential equations from (10), showing no systematic differences.

### v3

@@ Abstract
Cell tracking is a subject of active research gathering great interest in medicine and biology. Positron emission tomography (PET) is well suited for tracking radio-labeled cells in vivo due to its exceptional sensitivity and whole-body capability. For validation, ground-truth data are desirable that realistically mimic the flow of cells in a clinical situation. This study develops a workflow (CeFloPS) for simulating moving radio-labeled cells in a human phantom. From the XCAT phantom, the blood vessels are reduced to nodal networks along which cells can move and distribute to organs and tissues. The movement is directed by the blood flow, which is calculated in each node using the Hagen-Poiseuille equation and Kirchhoff’s laws assuming laminar flow. Organs are voxelized and movement of cells from artery entry to vein exit is generated via a biased 3D random walk. The probabilities of cells moving or remaining in tissues are derived from rate constants of tracer kinetic-based compartment modeling. PET listmode data is generated using the Monte-Carlo simulation framework GATE based on the definition of a large-body PET scanner with cell paths as moving radioactive sources and the XCAT phantom providing attenuation data. From the flow simulation of 100,000 cells, 100 sample cells were further processed by GATE and listmode data was reconstructed into images for comparison. As demonstrated by comparisons of simulated and reconstructed cell distributions, CeFloPS is capable of simulating cell behavior in whole-body PET. It achieves this simulation in a way that is anatomically and physiologically reasonable, thereby providing valuable data for the development and validation of cell tracking algorithms.
@@ 1 Introduction
Cell imaging and tracking are subjects of active research gathering great interest in the medical and biomedical science community. Non-invasive imaging techniques using fluorescent or luminescent markers in optical imaging (OI), iron oxide particles in magnetic resonance imaging (MRI), or radio-labeled tracers in single photon emission computed tomography (SPECT) and positron emission tomography (PET) have been frequently used in preclinical studies to permit in vivo visualization of living cells [1]. For human application, MRI, PET, and SPECT are the most promising tools as they provide either high spatial resolution (MRI) or excellent whole-body sensitivity (PET, SPECT). Of these imaging modalities, PET has become of special interest following labeled immune cells in the context of stem cell imaging and cell-based therapies [2].
PET is particularly advantageous because it combines exceptional sensitivity with the possibility of whole-body coverage, which has been further advanced in recent years by the design of large-field-of-view (LAFOV) PET systems with extended coverage of the axial field-of-view in the range of 1 to 2 meters, improved time-of-flight capabilities and excellent system sensitivities [3]. Using either long-lived isotopes or reporter genes, immunoPET has great potential to identify biomarkers allowing patients to be treated based on their most likely response to a dedicated immunotherapy [4, 5, 6].
The idea of imaging or tracking biological cells with PET is highly related to the field of Positron Emission Particle Tracking (PEPT) which has been developed primarily for technical applications, such as evaluation of particle–fluid or particle–particle interactions in the pharmaceutical, chemical, oil, mining, construction and power generation industries [7]. The idea is to follow one or a few radio-labeled particles by tracking the coincidence lines-of-response (LORs) recognized from its two back-to-back gamma rays emitted after beta-plus decay. In contrast to PET, where volumetric images of the radioactivity distribution are reconstructed from LORs, PEPT algorithms calculate a trajectory by connecting LORs in the most probable way, offering direct motion tracking of radio-label particles [8]. As PEPT applications are generally carried out outside a clinical environment, special industrially usable devices were developed in parallel to the invention of clinical PET systems [9, 10]. Following the idea of PEPT in a clinical or pre-clinical environment, the outstanding sensitivity of PET can be utilized to track radio-labeled singular cells traveling inside a biological system (e.g. human, mouse) [10, 11, 12, 13].
@@
The second simulation enables the generation of raw PET data from the simulated cell pathways using Monte Carlo based physics simulations. Such simulations were also used in the PEPT context based on the Geant4 Application for Tomographic Emission (Gate) [19]. For example, Herald et al. simulated both the ADAC/Phillips Forte and the Siemens Inveon pre-clinical PET and compared fluid simulations with experimental data [20, 21]. The foundation of our Gate simulation, developed here, is the concept of new total-body PET with an extensive field of view and high sensitivity, which presents novel possibilities for human whole-body cell tracking.
The task of cell tracking is not part of this work, but has been developed in many variants by different groups in the past [22, 5]. Nguyen et al. recently presented a new algorithm, PEPT-EM, which was able to successfully track the fate of more than 70 radio-labeled melanoma cells in a mouse model [13]. In this context, the group recently introduced an image reconstruction algorithm for PET that employs an optimal transport-based regularisation technique. This algorithm demonstrates considerable potential for the tracking of radioactively labeled cells within a human body, with no limitation on the number of trackable cells, and with the capacity to address the inherent handling of scattered, random, and background events. The proposed methodology integrates the information from all detected events to facilitate the reconstruction of dynamic evolution, while concurrently enhancing the reconstruction quality through the enforcement of temporal consistency. [23, 24, 25]
@@ 2.1 Simulation basis
CeFloPS is based on anatomical data provided by the XCAT phantom with high-resolution 3D models of all organs, blood vessels, bones, and muscles (Fig. 1a) in the form of STEP files [26]. Using Inventor (Version 2023, Autodesk Inc., San Rafael, USA), these STEP files were converted into StL files (StereoLithography file format), yielding triangulated surface meshes of organs and structures. The StL files were further processed in Python (Version 3.11.5, Python Software Foundation) with the help of the trimesh library [27], allowing for operations on the meshes and giving access to the coordinates of the vertices. For further processing, meshes of muscles and internal organs were voxelized at a resolution of 0.5 mm $\times$ 0.5 mm $\times$ 0.5 mm, whereas the complex anatomy of blood vessel meshes was simplified by reducing it to node networks. The nodes describe the possible paths a cell can take and are placed along the central axis of a blood vessel, storing information about (i) its spatial coordinates, (ii) which nodes are adjacent in the direction of blood flow and (iii) the radius of the blood vessel.
To generate the nodes, the blood vessels were divided into segments starting from one end point of a vessel. For each segment, the barycenter was formed of all vertices located in this segment, resulting in points lying inside a blood vessel. Starting from the first point inside a blood vessel, planes, orthogonal to the blood vessel, were then placed at an interval of 1 mm along the vectors between these points. Using the intersection of those planes with the blood vessel surface, the vessel radius and the new center point were calculated, resulting in the final node positions spaced at 1 mm (see Fig. 2).
Nodes that lie at bifurcations of blood vessels and thus have more than one neighbor in the direction of blood flow were marked and are hereafter referred to as links. Since the spatial resolution of the XCAT phantom is limited, some connections between vessels are missing, as seen in Fig. 1b. In such cases, we interpolated linearly between the nodes of those vessels to generate a connection. In addition, the spatial resolution of blood vessels varies from organ to organ. The smallest radius defined in XCAT is 0.15 mm and is located in the toes. Consequently, no capillary bed is included that connects the nodes of arteries and veins. (Fig. 1c). This results in a total of five separate networks of nodes: arteries, veins, portal vein, and pulmonary arteries and veins. All of these, except the portal vein, are connected via the heart. In order to compensate for the lack of capillaries, these separate networks are artificially combined to form a single closed network in which opposing arteries and veins leading to the same organ are virtually connected.
@@ 2.2 Blood flow simulation
From the node networks, the mean blood flow rate $\overline{Q}_{i}$ and the vessel resistance $R_{i}$ were estimated for each node $i$ from a highly simplified flow model using Kirchhoff’s laws in combination with the Hagen-Poiseuille equation. This was done by assuming laminar flow in every vessel, each treated as a cylinder [28, 29]:
@@
$\mathrm{Re}=\frac{2\rho\overline{Q}_{i}}{\pi r_{i}\eta}\,$, (5)
with $\rho$ the density of blood set to 1.06 g/ml. Laminar flow is assumed valid if the Reynolds number is below 2000, which is the known threshold beyond which laminar flow gradually transitions into turbulent flow (Reynolds number > 4000) [30]. Due to the lack of small blood vessels in the XCAT phantom, our flow model misses a large portion of high resistances which would add to the calculation of the total resistance $R_{\mathrm{tot}}$ in (3) and (4). By (1), this also affects the blood flow $\overline{Q}_{i}$ in each node.
At a link, the probability $p_{\mathrm{node}}$ with which a cell will move to the next possible node is calculated based on the blood flow $\overline{Q}_{\mathrm{link}}$ in the link and $\overline{Q}_{i}$ in each following node:
@@
The velocity of an individual cell is assigned by pseudo-random number sampling based on a cumulative distribution function (CDF). This ensures that a group of cells starting from the same node have different velocities and disperse over time in the blood vessels. The CDF was derived using the velocity profile of laminar flow in a tube [31]
@@ 2.3 Organ distribution simulation
Inside a voxelized organ, cell motion is guided by a biased 3D random walk along the voxels to simulate the capillary bed and the tissue exchange. This process is carried out in a series of discrete calculations, each with a defined time step. The current simulations have set the time step to 600 ms. First, voxels containing an artery entrance or a vein exit were identified and hereafter referred to as start-voxels and end-voxels, respectively. If an organ does not have an artery entrance and/or vein exit due to missing connections in the XCAT phantom, the nearest artery/vein is calculated and the closest voxel to that blood vessel is determined as start-/end-voxel.
Remaining blood vessels are connected to an organ in the same way if their distance to that organ lies within a 2 cm threshold. To compensate for missing arterial branches in muscles and organs, each node of the arteries and veins can serve as entrance and exit for cells, respectively. For every voxel within an organ or muscle, direction vectors towards all start- and end-voxels were determined along with separate weighting values. As the distance $d$ between a voxel and its start/end voxel increases, these weightings decrease by a factor of $1/d^{2}$, as shown in Fig. 3. The vectors point away from start-voxel or towards end-voxel. The combination of those vectors represents the most probable directions for a cell to move. The weightings of vectors in start- and end-voxels were set to zero.
The random walk is initiated at a start-voxel, where a cell can move to one of the surrounding voxels. This yields 26 possible directions (provided the start-voxel does not reside on the organ’s periphery), all with the same probability. For the proceeding steps, the direction is biased in accordance with the previously determined weightings of the direction vectors in each voxel. If a cell reaches an end-voxel, the random walk concludes, followed by the cell exiting via the vein back into the nodal network for further blood flow simulation. The speed at which cells perform the random walk in the different organs was calibrated against the delay times determined by [32]. For this purpose, first the distribution of required time to travel from start- to end-voxel in each organ is determined for cells at an arbitrary random walk speed. The mean value of this distribution is calculated and compared with the value derived by [32], leading to new speed values for the corresponding organ.
@@
In addition to the biased random walk, a compartment model was incorporated into the organ distribution simulation to account for the transition state of a cell in tissue. Besides its spatial voxel location, each cell occupies a compartmental state, which dictates whether a cell can move or not. The exchange between compartments is determined by transition rates $k_{i}$. The number of compartments and the values of the transition rates must be known in advance to the CeFloPS simulation and can be obtained from (physiologically based) kinetic models. As a first example and as proof-of-principle, we employ a two-compartment model comprising the compartments $C_{1}$ and $C_{2}$ (shown in Fig. 4). The compartment $C_{1}$ signifies a reversible binding of the cells, whereas $C_{2}$ denotes an irreversible binding. Consequently, if a cell is situated within either $C_{1}$ or $C_{2}$, it does not execute a random walk and remains temporarily or permanently in the same voxel. The change in amount of cells in these compartments can be described with the transition rates $K_{1}$, $k_{2}$, and $k_{3}$ through the differential equations
@@
The arterial input function (AIF), denoted by $C_{A}(t)$, serves as a metric for the concentration or number of cells present in the arterial blood and also has to be entered as input into the CeFloPS simulation. One method for generating an AIF is outlined in [33]. It is important to note that, in general, transition rates are only defined in the underlying kinetic model for a subset of the tissues included in CeFloPS. Consequently, the sum of all cells transitioning from the arterial blood into the specified tissues may not be equal to the rate of change of the AIF. In CeFlops, this would result in an erroneous distribution of cells, which would not correspond to (11). To counteract this, two virtual compartments were introduced. The first compartment, denoted by $C_{\mathrm{inj}}$, represents the injection of cells into the arterial blood during the beginning of the simulation. The second compartment, designated by $C_{\mathrm{rem}}$, describes cells which are removed from the arterial blood and go into a tissue not defined in the kinetic model rates. The rate of change of the AIF is then given by
@@
The transition rates $k_{\mathrm{inj}}$ and $k_{\mathrm{rem}}$ are selected as time-dependent to account for the time-limited injection phase of the cells at the beginning of a simulation and potential delays during the transition into tissues. The determination of $k_{\mathrm{inj}}(t)$ and $k_{\mathrm{rem}}(t)$ is achieved through the implementation of a 3-layer neural network. The neural network identifies the optimal values for $k_{\mathrm{inj}}$ and $k_{\mathrm{rem}}$ at each simulation time $t$ by fitting the cell distributions in all tissues to fulfill the differential equations from (11) and (12). The latter phase of the distributions, in which a state of equilibrium is ultimately achieved, is given greater weight relative to the initial phase, during which significant alterations may occur. This ensures that the equilibrium phase at late stages of the simulation is primarily governed by the compartment model. The neural network is set up using the Python library PyTorch [34].
@@
In this example, we implement the rate constants of a 18F-Fluordesoxyglucose (FDG) model taken from [39, 40] and listed in Table 1. The application of kinetic modeling with FDG has been comprehensively addressed in the extant literature, thereby facilitating a more robust validation of our simulation compared to kinetic modeling with cells, which is still scarce in the literature and usually only takes few tissues into account. The simulated cells in this example therefore do not behave like real, existing cells, but like an FDG tracer. In this context, $C_{1}$ and $C_{2}$ denote the unbound and bound states of an FDG ”cell” in tissue and $C_{A}$ the free FDG ”cell” in the blood supply of the tissue. Based on the probabilities calculated with (13) from these transition rates, a random decision is made at each time step as to whether a cell transitions between these compartments. Fig. 5 demonstrates the temporal cell distributions calculated by the Markov chain approach in comparison to the whole population based compartment modelling using differential equations from (11), showing no systematic differences.